# MIM-Refiner: A Contrastive Learning Boost from Intermediate Pre-Trained Representations

### v2

@@ E.3 ImageNet-1K Low-shot Evaluation Details
E.4 ImageNet-1K $k$-NN Classification Details
@@
E.5 ImageNet-1K Linear Probing Evaluation Details
@@
As the linear probes trained on the concatenation of the last 4 blocks have more features and more parameters to discriminate between classes, they tend to be the best within the sweeped parameters. Note that we evaluate the representation of the last block in isolation via $k$-NN classification. We investigate $k$-NN classification with features from the last 4 blocks in Appendix C.11.
E.6 ImageNet-1K Cluster Evaluation Details
@@
E.7 iNat18 Transfer Learning Evaluation Details
@@
E.8 Transfer Learning Linear Probing
For transfering the pre-trained features to iNat18, six VTAB datasets and ADE20K (Table 4) we use the DINOv2 [56] linear probing protocol as described in Appendix E.5. For ADE20K and iNat18, we reduce the hyperparameter grid to fit into 40GB GPU memory.
E.9 ADE20K Semantic Segmentation Linear Probe
@@
E.10 Fine-tuning on VTAB-1K
@@
E.11 Fine-tuning With 100% Labels
@@
E.12 ADE20K Semantic Segmentation Fine-tuning
We fine-tune ViT-L models using a linear head to predict a class label per 16x16 patch which is then upscaled to the original resolution. We follow common practices and fine-tune on 512x512 resolution, where we interpolate the absolute positional embedding from 224x224 to 512x512, add relative position bias to the attention layers (initialized to 0) [29] and introduce layerscale [67] (initialized to 1). A common augmentation pipeline is used that consists of random rescaling, random horizontal flipping, color jitter and padding if necessary. We train for 40K updates using a batchsize of 16, a learning rate of 1e-4, weight decay 0.05, linear warmup for 1.5K update steps followed by cosine decay, stochastic depth rate 0.2, layer-wise learning rate decay 0.95. We evaluate after 40K update steps using a sliding window of 341 pixels and report mIoU over the validation set.
This setting closely follows [41], who showed that the performance of this relatively simple pipeline has a high correlation to performance larger models and more sophisticated segmentation pipelines.
E.13 Loss Schedule Visualizatios
We visualize the schedules used for scheduling the loss weight of ID heads attached at intermediate blocks in the ablation study (Table 4(f)) in Figure 9.
Figure 9: Loss weight schedules for the ablation in Table 4(f). For visual clarity, small offsets are added when values overlap (for “One Hot” and “Staggered Step”).

### v4

@@ D.3 ImageNet-1K Low-shot Evaluation Details
D.4 ImageNet-1K $k$-NN Classification Details
@@
D.5 ImageNet-1K Linear Probing Evaluation Details
@@
As the linear probes trained on the concatenation of the last 4 blocks have more features and more parameters to discriminate between classes, they tend to be the best within the sweeped parameters. Note that we evaluate the representation of the last block in isolation via $k$-NN classification. We investigate $k$-NN classification with features from the last 4 blocks in Appendix B.15.
D.6 ImageNet-1K Cluster Evaluation Details
@@
D.7 iNat18 Transfer Learning Evaluation Details
@@
D.8 Transfer Learning Linear Probing
For transfering the pre-trained features to iNat18, six VTAB datasets and ADE20K (Table 4) we use the DINOv2 [62] linear probing protocol as described in Appendix D.5. For ADE20K and iNat18, we reduce the hyperparameter grid to fit into 40GB GPU memory.
D.9 ADE20K Semantic Segmentation Linear Probe
@@
D.10 Fine-tuning on VTAB-1K
@@
Table 27: Hyperparameters for fine-tuning on VTAB-1K.
D.11 Fine-tuning With 100% Labels
@@
D.12 ADE20K Semantic Segmentation Fine-tuning
We fine-tune ViT-L models using an UperNet [89] segmentation head to predict a segmentation mask. We follow common practices and fine-tune on 512x512 resolution, where we interpolate the absolute positional embedding from 224x224 to 512x512, add relative position bias to the attention layers (initialized to 0) [31] and introduce layerscale [77] (initialized to 1). A common augmentation pipeline is used that consists of random rescaling, random horizontal flipping, color jitter and padding if necessary. We train for 160K updates using a batchsize of 16, a learning rate of 2e-5, weight decay 0.05, linear warmup for 1.5K update steps followed by cosine decay, stochastic depth rate 0.2, dropout 0.1, layer-wise learning rate decay 0.95. We evaluate after 160K update steps using a sliding window of 341 pixels and report mIoU over the validation set.
D.13 $k$-NN Classification and Semantic Segmentation Probe per Block
For the per-block analysis in Figure 3(b) and Figure 8 we follow the respective settings of $k$-NN classification (Appendix Section D.4) and semantic segmentation linear probing (Appendix Section D.9). The only change is that for semantic segmentation linear probing per-block, we fix the learning rate to 0.1 and only use the patch tokens of the respective block as input to the linear probe.
D.14 Loss Schedule Visualizatios
We visualize the schedules used for scheduling the loss weight of ID heads attached at intermediate blocks in the ablation study (Table 1) in Figure 13.
Figure 13: Loss weight schedules for the ablation in Table 1. For visual clarity, small offsets are added when values overlap (for “One Hot” and “Staggered Step”).
Appendix E Representation Degradation in Other Modalities
We investigate feature degradation of masked pre-trained models from different domains in Figure 14 which show similar trends, suggesting that our methodology could be extended to other modalities.
(a) VideoMAE $\rightarrow$ ImageNet-1K
(b) RoBERTa $\rightarrow$ MNLI
(c) AudioMAE $\rightarrow$ SpeechCommands V2
(d) AudioMAE $\rightarrow$ ESC-50
Figure 14: Feature degradation investigation on other modalities. (a) We evaluate VideoMAE [76] on ImageNet-1K classification by treating each image as a single frame video. (b) Entailment classification using RoBERTa [56] on the MNLI [86] task of the GLUE [81] benchmark. (c) AudioMAE [40] is evaluated on SpeechCommandsV2 [84] audio classification (d) AudioMAE [40] is evaluated on ESC-50 [66] environmental sound classification. Similar trends can be observed in other modalities, often even on smaller models.